OpenStudy (annie12m):

I medal if you check my work. Kendall has a maximum of 500 megabytes of data usage each month on her cell phone. She has used 160 megabytes this month. Each message she sends is 4 megabytes. About how many more messages can Kendall send this month? A. at least 85 messages B. at most 85 messages C. at least 165 messages D. at most 165 messages 500 divided 4 = 125 messages total each month 160 divided 4 = 40 messages sent 125 - 40 = 85 messages left so the answer is B?

OpenStudy (annie12m):

@aleroth @b87lar @CheyenneCoffelt @laila22 @surjithayer

OpenStudy (mathmale):

Thank you for starting your own work on this problem and for sharing that work. Note that Kendall has already used some of his megabyte allotment, so you should not be starting with 500 mb. Instead, figure out how many mb are still available to this guy.

OpenStudy (annie12m):

500 - 160 = 340 mb left

OpenStudy (annie12m):

i actually did this first, but there wasnt an answer when i finished the work

OpenStudy (anonymous):

\[160+4x \le 500,\]

OpenStudy (annie12m):

waittt i know what i did wrong the first time i tried dividing only 34 tthen adding the zero to the end thats why it didnt work

OpenStudy (annie12m):

340 divided 4 equals 85

OpenStudy (cheyennecoffelt):

yes, so how many texts can she send? 340/4=?

OpenStudy (annie12m):

both ways equal same answer

OpenStudy (cheyennecoffelt):

okay, can she send more than the quotient?

OpenStudy (annie12m):

no thats the most she can send

OpenStudy (annie12m):

so still b

OpenStudy (cheyennecoffelt):

So now we know it's not A. yeah

OpenStudy (anonymous):

B is the correct solution.

OpenStudy (cheyennecoffelt):

160+4(85)≤500,

OpenStudy (anonymous):

\[4x \le 500-160\] or \[4x \le340\] or \[x \le \frac{ 340 }{ 4 }\] or \[x \le85\]

OpenStudy (anonymous):

or at the most 85 messages.